OpenStudy (anonymous):

ANSWER QUICKLY!!! MEDAL, BEST RESPONSE AND FAN RIGHT AWAY!!! what is the slope of the line defined by the equation 8x + 2y = 5

OpenStudy (anonymous):

-8/2

OpenStudy (anonymous):

m=-4 b=5/2 where m is the slope and b is the y intercept

OpenStudy (anonymous):

8x+2y=5 subtract 8x from both sides 2y=-8x+5 now divide both sides by 2 y=-4+5/2 m=-4